I need help with these problems. Could you please help me? I do them over and over and they dont make sense when I check them. (they are inequalities
1) -(6b-2)>O
2)-2(r-3)+7 (is greater than or equal to) 8
Thanks

Answer by stanbon(75887)

1) -(6b-2)>O
-------------
Note: When an inequality is multiplied or divided on both sides by a
negative, the inequality sign is reversed.
-------------
If you multiply both sides of your problem be (-1) you get:
6b-2 < 0
6b < 2
b < 1/3
---------------------
2)-2(r-3)+7 >= 8
-2r + 6 + 7 >= 8
-2r + 13 >=8
-2r > = -5
2r <= 5
x <= 5/2
==================
Cheers,
Stan H.